OpenStudy (waheguru):

−11 + 10( p + 10) = 4 − 5(2p + 11) Why am I gettting 7 rather than -7?

OpenStudy (waheguru):

Here is How I solved it...

OpenStudy (anonymous):

-11+10(p+10)=4-5(2p+11) -11+10p+100=4-10p-55 10p+10p=4-55+11-100 20p=-140 p=-7

OpenStudy (waheguru):

−11 + 10( p + 10) = 4 − 5(2p + 11 -11+10p+100=4-10p-55 10p+89=-51-10p 89=-51-10p-10p 89=-51-20p 89+51=-20p 140=-20p 140/20 = p 7 =p?

OpenStudy (anonymous):

140=-20p 140/20 = p <--- mistake appears. a negative sign is missing

OpenStudy (waheguru):

WHen we bring it over dont we remove the negative sign

OpenStudy (waheguru):

Because i Houhgt it becomes opposite operation

OpenStudy (anonymous):

we will never remove negative sign. To make it positive, we have to multiply (-1) to both sides

OpenStudy (anonymous):

140=-20p 140/20=-p (-1)(140/20)=(-1)(-p) -7=p

OpenStudy (waheguru):

10p+89=-51-10p 89=-51-10p-10p here when I brought the 10p over I changed it to -10p why do I do the same to the -140 and make It positive when I bring it over

OpenStudy (anonymous):

10p+89=-51-10p 10p+89-10p=-51-10p-10p <--- minus 10p to both sides 89=-51-20p

OpenStudy (waheguru):

I think I got it thanks

OpenStudy (anonymous):

you are welcome